OpenStudy (anonymous):

D is the midpoint of line CE. E has coordinates (-3,5), and D has coordinates (2 1/2,1). Find the coordinates of C.

OpenStudy (anonymous):

C is (8,-3)

OpenStudy (anonymous):

let C to be (x,y) then solve for x in [ x+(-3)]/2=2 1/2 , and solve for y in (y+5)/2 =2

OpenStudy (anonymous):

\[(\frac{x_C+x_E}{2},\frac{y_C+y_E}{2}) = (x_D+y_D)\] \[E = (-3,5) = (x_E, y_E)\]\[D=(2\frac{1}{2},1) = (x_D,y_D)\] \[(\frac{x_C+(-3)}{2},\frac{y_C+5}{2}) = (2\frac{1}{2}+1)\] \[\frac{x_C-3}{2}=2\frac{1}{2}=\frac{5}{2}\]\[\frac{y_c+5}{2}=1\] \[x_C-3=5 => x_C=8\]\[y_C+5=2 => y_C=-3\]